Q3: Filer Manufacturing has II million shares of common stock outstanding. The current share price is $63 and the book value per share is S6+ Filer Manufacturing also has two bond issues outstanding. The fi rst bond issue has a face value of $70 million, has a 7 percent coupon, and sells for 93 percent of par. The second issue has a face value of $55 million, has an 8 percent coupon, and sells for 104 percent of par. The f rst issue matures in 21 years, the second in 6 years. a. What are Filers capital structure weights on a book value basis? b. What are Filers capital structure weights on a market value basis? c. Which are more relevant, the book or market value weights? Why? 2
Answer #1

a)Book Value of Equity= $6 x 11,000,000 = $66,000,000

Book Value of Debt= $70,000,000+ $55,000,000 = $ 125,000,000

Book Value of Filer = $66,000,000+ $125,000,000

=$191,000,000

The book value weights of equity and debt are:

Equity= 66,000,000/191,000,000= 0.34554974

Debt = 125,000,000/191,000,000= 0.65445026

b)

Market Value of Equity= $68 x 11,000,000 = $748,000,000

Market Value of Debt= $70,000,000 x 0.93+ $55,000,000 x 1.04 = $ 122,300,000

Market Value of Filer = $748,000,000+ $122,300,000

=$870,300,000

The Market value weights of equity and debt are:

Equity= 748,000,000/870,300,000= 0.85947374

Debt = 122,300,000/870,300,000= 0.14052626

C) The Market Value are more Relevant as its Represents the market's Expectations of future cash flows discounted however book value is historic price